OpenStudy (anonymous):

prove the following identities cos(x-pi) = sin(x+3pi/2) tan(2x) = 2tan(x)/1-tan^2(x) (hint switch to sine and cosine)

OpenStudy (matlee):

why is yours red

OpenStudy (anonymous):

use these identities for the 1st one \[\cos(A-B)=cosAcosB+sinAsinB\] \[\sin(A+B)=sinAcosB+sinBcosA\]

OpenStudy (matlee):

Qualified helpers dont come , i just bought some owl bucks for them and they didnt com for 30 minutes and still ahvent come anyways. I ased for a qualified helper too

OpenStudy (anonymous):

for that second one use this identity-\[\tan(A+B)=\frac{ tanA+tanB }{ 1-tanAtanB}\] try putting A=B=x in this question :)

OpenStudy (anonymous):

@lisamath12

OpenStudy (anonymous):

ok so for the second one that leaves me with tan(x+x) = tanx+tanx/1-(tanx)(tanx)

OpenStudy (anonymous):

I'm not sure where to go from there though

OpenStudy (anonymous):

yes :) ur correct now try to simplify further

OpenStudy (anonymous):

I am not sure how to do that I have all the formulas for these questions what I don't know is how to solve it after that

TheSmartOne (thesmartone):

Do we have to prove one or two identities for this question?

OpenStudy (triciaal):

cos(A-B) = Cos (A + B) + 2 Sin A Sin B

TheSmartOne (thesmartone):

For the second identity, \(\sf\Large tan(2x) = \frac{sin(2x)}{cos(2x)}\)

TheSmartOne (thesmartone):

Another hint for the second one: \(\sf\Large sin(2x) =2sin(x)cos(x)\) and \(\sf\Large cos(2x) =cos^2(x) -sin^2(x)\)

TheSmartOne (thesmartone):

@lisamath12 You will get your answer faster if you cooperate :)

OpenStudy (anonymous):

sorry I was busy for a short while just got back

TheSmartOne (thesmartone):

ok :)

OpenStudy (anonymous):

so now we have: 2sin(x) cos(x)/ Cos^2(x) - sin^2(x) is equal too 2tan(x)/1-tan^2(x)

OpenStudy (anonymous):

is the next step to change the tangents on the other side too sine and cosine as well?

TheSmartOne (thesmartone):

we only deal with one side at a time :)

TheSmartOne (thesmartone):

so we have \(\sf\Large \frac{2sin(x)cos(x)}{cos^2(x) -sin^2(x)}\) divide both the numerator and denominator by cos^2(x)

TheSmartOne (thesmartone):

\(\sf\Huge \frac{\frac{2sin(x)cos(x)}{cos^2(x)}}{\frac{cos^2(x) -sin^2(x)}{cos^2(x)}}=??\)

TheSmartOne (thesmartone):

So lets deal with the numerator first :) \(\sf\Large \frac{2sin(x)cos(x)}{cos^2(x)} = ?\) Simplify it :)

OpenStudy (anonymous):

alright so now we got 2sin(x)/cos(x)

OpenStudy (anonymous):

and the denominator is -sin^2(x)?

TheSmartOne (thesmartone):

simplify 2sin(x)/cos(x) hint: tan(x) = sin(x)/cos(x)

OpenStudy (anonymous):

2 tan(x)

TheSmartOne (thesmartone):

exactly! :D

TheSmartOne (thesmartone):

ok, now we simplify the denominator part :)

TheSmartOne (thesmartone):

\(\sf\Huge \frac{cos^2(x) -sin^2(x)}{cos^2(x)}=??\)

TheSmartOne (thesmartone):

Hint: (a-b)/c = a/c - b/c

OpenStudy (anonymous):

-sin^2(x)/cos^2(x)

TheSmartOne (thesmartone):

Not sure how you got that, but :P \(\sf\Large \frac{cos^2(x) -sin^2(x)}{cos^2(x)}=\frac{cos^2(x)}{cos^2(x)} - \frac{sin^2(x)}{cos^2(x)} = ?\)

OpenStudy (anonymous):

oh is it - tan(x). I had cos^2(x)/cos^2(x) equal 0 and that lefted me with -sin^2(x)/cos^2(x)

TheSmartOne (thesmartone):

a/a = 1 \(\sf\Large \frac{cos^2(x)}{cos^2(x)} = 1\) so we are left with 1 - tan^2(x)

TheSmartOne (thesmartone):

And therefore we have proved the identity

TheSmartOne (thesmartone):

\(\sf\Huge \frac{\frac{2sin(x)cos(x)}{cos^2(x)}}{\frac{cos^2(x) -sin^2(x)}{cos^2(x)}}=\frac{2tan(x)}{1-tan^2(x)}\)

OpenStudy (anonymous):

ohhh, yes I see it now thank you!

TheSmartOne (thesmartone):

\(\sf\Large tan(2x) = \frac{2tan(x)}{1-tan^2(x)}\) LHS = RHS

TheSmartOne (thesmartone):

Anytime! :D

TheSmartOne (thesmartone):

Now we have this question, right? cos(x-pi) = sin(x+3pi/2)

OpenStudy (anonymous):

yes exactly

TheSmartOne (thesmartone):

Hold on, I'm trying to see how to solve this one. :)

TheSmartOne (thesmartone):

So, I got it now :)

TheSmartOne (thesmartone):

we need to use these two formulas cos(α – β) = cos(α)cos(β) + sin(α)sin(β) sin(α + β) = sin(α)cos(β) + cos(α)sin(β)

TheSmartOne (thesmartone):

We shall seperately open up the LHS and RHS :)

TheSmartOne (thesmartone):

Or we could use this actually: cos(α – β) = cos(α + β) + 2sin(α)sin(β)

TheSmartOne (thesmartone):

So @lisamath12 lets plug it into the last formula I gave :)

OpenStudy (anonymous):

alright so then we would have: cos(x-pi) = cos(x+pi) + 2sin(x)sin(pi)

TheSmartOne (thesmartone):

and lets first expand cos(x+pi) using this formula: cos(α + β) = cos(α)cos(β) – sin(α)sin(β)

OpenStudy (anonymous):

cos(x+pi) = cos(x)cos(pi) - sin(x)sin(pi)

TheSmartOne (thesmartone):

lets simplify that hint: cos(pi) = -1 sin(pi) = 0

OpenStudy (anonymous):

alright so cos(x+pi) = cos(x)(1) - sin(x)(0)

TheSmartOne (thesmartone):

cos(x)(-1) ** so simplifying that gets us -cos(x) right?

TheSmartOne (thesmartone):

and let's simplify the other half 2sin(x)sin(pi)

TheSmartOne (thesmartone):

remember sin(pi) = 0

OpenStudy (anonymous):

2sin(x)(0)=0

TheSmartOne (thesmartone):

correct so we got the LHS as -cos(x) lets expand the RHS too

TheSmartOne (thesmartone):

sin(x+3pi/2) using the formula: sin(α + β) = sin(α)cos(β) + cos(α)sin(β)

OpenStudy (anonymous):

sin(x+3pi/2) = sin(x)cos(3pi/2) + cos(x)sin(3pi/2)

OpenStudy (anonymous):

sin(x)(0) + cos(x)(-1)

TheSmartOne (thesmartone):

and simplifying it more you get -cos(x)

TheSmartOne (thesmartone):

therefore -cos(x) = -cos(x) and LHS = RHS

OpenStudy (anonymous):

once again big help

TheSmartOne (thesmartone):

anytime :)